### v1

# Agentic Multi-Persona Framework for Evidence-Aware Fake News Detection

Abstract.
The rapid proliferation of online misinformation poses significant risks to public trust, policy, and safety, necessitating reliable automated fake news detection. Existing methods often struggle with multimodal content, domain generalization, and explainability. We propose AMPEND-LS, an agentic multi-persona evidence-grounded framework with LLM–SLM synergy for multimodal fake news detection. AMPEND-LS integrates textual, visual, and contextual signals through a structured reasoning pipeline powered by LLMs, augmented with reverse image search, knowledge graph paths, and persuasion strategy analysis. To improve reliability, we introduce a credibility fusion mechanism combining semantic similarity, domain trustworthiness, and temporal context, and a complementary SLM classifier to mitigate LLM uncertainty and hallucinations. Extensive experiments across three benchmark datasets demonstrate that AMPEND-LS consistently outperformed state-of-the-art baselines in accuracy, F1 score, and robustness. Qualitative case studies further highlight its transparent reasoning and resilience against evolving misinformation. This work advances the development of adaptive, explainable, and evidence-aware systems for safeguarding online information integrity.
Fake News Detection, Misinformation Detection, Information Retrieval, LLM
$c$s: Computing methodologies Reasoning about belief and knowledgeccs: Information systems Information retrieval
1. Introduction
News plays a central role in shaping public opinion, guiding decisions, and maintaining social stability. With digital media and social platforms accelerating global dissemination, information now spreads almost instantly. However, this connectivity also amplifies the reach of fake news (zhou2020survey), which can erode trust, cause confusion, and trigger severe consequences in health, politics, and public safety. These societal risks have made fake news detection a critical research area for safeguarding the integrity of online ecosystems (shu2020fakenewsnet; das2025fakenewsdetectionllm). Early work on fake news detection relied on manual fact-checking, reliable but unscalable (graves2018understanding). Automated approaches reframed detection as supervised classification, leveraging linguistic cues, user behavior, and propagation patterns (shu2017fake). Deep learning further advanced these models by capturing complex textual, semantic, and contextual dependencies (zhou2020survey). Yet, the rapid evolution of misinformation and heterogeneous news sources reveals persistent gaps, particularly in multimodal integration, cross-domain generalization, and reliance on resource-intensive annotations.
Over time, methods have evolved from text-only classifiers to multimodal, domain-adaptive, and large language model (LLM)-driven paradigms. Multimodal approaches, from early joint encoders EANN (wang2018eann)/ MVAE (khattar2019mvae) to contrastive and graph-enhanced models (wei2025structure; 10.1145/3581783.3612423), demonstrate the value of heterogeneous cues, but suffer from fixed encoders, shallow fusion, and costly pipelines. Domain adaptation methods (WU2025111485; 10.1145/3664647.3681317) mitigate distribution shifts yet depend on domain metadata and remain brittle under unseen domains. More recently, LLM-based methods (hu2025synergizingllmsgloballabel; modzelewski2025pcotpersuasionaugmentedchainthought) enhance reasoning and semantic understanding, but face high computational costs, weak multimodal integration, and shallow, non-evidence-grounded explanations, limiting use in high-stakes settings. Collectively, existing approaches capture correlations but lack scalability, robustness, and evidence-grounding transparency (harris2024review).
To address these challenges, we propose an explainable, evidence-grounded multi-persona agentic framework for multimodal fake news detection. Our framework integrates evidence retrieval through text-based web search and reverse image search, with each piece of evidence evaluated by a composite reliability score combining lexical and semantic similarity, domain credibility, and temporal consistency. In parallel, a knowledge graph (KG) encodes entity–relationship structures to provide factual grounding. These heterogeneous signals are synthesized by a multi-persona agentic module that engages in multi-round, persona-driven questioning to iteratively refine its contextual memory. Building on this, we design a modular classification pipeline where an LLM-based classifier produces pseudo labels with justifications, and a lightweight small language model (SLM) distills both the agent’s reasoning and LLM outputs. Importantly, for uncertain cases, we add persuasion-based analysis that detects manipulative strategies, computes a persuasion index, and refines reasoning. This design ensures low latency, high efficiency, and strong explainability, advancing multimodal fake news detection.
Our key contributions are summarized below:
(i) Complementary Agentic Reasoning with Persuasion Refinement and SLM Integration: We develop an LLM-based multi-persona reasoning framework where specialized agents (supervisor, journalist, legal, scientific) iteratively interrogate evidence. Uncertain cases are re-evaluated using a persuasion-knowledge module that detects manipulative strategies. The distilled reasoning is transferred to a lightweight SLM classifier, making the pipeline both interpretable and computationally efficient.
(ii) Cross-Modal Evidence Retrieval with Reliability Fusion: To support agentic reasoning, we design a robust evidence retrieval pipeline spanning text, images, and knowledge graphs. Candidate documents are ranked using a hybrid reliability score that fuses semantic similarity, domain credibility, and temporal consistency, ensuring complementary signals from multiple modalities reinforce factual grounding.
(iii) Unified Evaluation of Complementary Strategies: Experiments on benchmark datasets and real-world case studies, we show that the combination of agentic reasoning, persuasion refinement, and reliability-guided multimodal evidence provides consistent improvements over individual baselines. Our component-wise analysis further indicates that each strategy contributes in complementary ways toward enhancing accuracy, efficiency, and interoperability.
2. Related Work
Fake news detection is a challenging classification task requiring reasoning over diverse information sources. Early methods struggled with multimodal evidence, while recent baselines integrate text and images but rely on simple fusion. Current advances can be broadly grouped into three categories: multimodal techniques (khattar2019mvae; wei2025structure), domain adaptation methods, and advanced language model–based approaches.
(i) Multimodal Fake News Analysis: Early multimodal methods, such as EANN (wang2018eann) and MVAE (khattar2019mvae), combined textual and visual features but relied on fixed encoders with limited adaptability. Subsequent works tackled cross-modal inconsistencies via similarity- and ambiguity-aware fusion (SAFE (zhou2020similarity), CAFE (chen2022cross)), while SpotFake (singhal2019spotfake) integrated BERT for text yet retained shallow fusion. More recently, contrastive learning has dominated, with StruACL (wei2025structure), COOLANT (10.1145/3581783.3613850), and MFCL (CHEN2025112800) improving alignment through adversarial or contrastive objectives, and GAMED (10.1145/3701551.3703541) exploiting mixture-of-experts for feature discrimination. Graph-based approaches enhance this by incorporating semantic and knowledge graphs (HSEN (10.1145/3581783.3612423), KAMP (10.1145/3696410.3714532), GS2F (10.1145/3708536)). Transformer-based architectures further refine fusion with hierarchical attention (10.1145/3581783.3612423; 10.1145/3581783.3613850; 10.1145/3696410.3714532; li-etal-2025-imrrf; 10834537). Despite these advances, most methods still rely on pretrained encoders, incur high computational cost, and lack explicit evidence-grounded reasoning, explainability, and scalability, limiting robustness in high-stakes misinformation detection.
Figure 1. Proposed Framework: AMPEND-LS
(ii) Domain Adaptation-based Methods: Domain adaptation is extensively studied to handle distribution shifts in fake news detection (wei-etal-2025-cross). Recent methods include hierarchical and transformer-based architectures such as HFGD (WU2025111485), MMDFND (10.1145/3664647.3681317), BREAK (10834537), and StruACL-TGN (wei2025structure), which leverage attention or structure-aware mechanisms for cross- and multi-domain settings. Transfer learning approaches like MMHT (10.1145/3696410.3714517) disentangle veracity-relevant from domain-specific features, while PFND (11027119) applies graph convolutional networks to combine content and fact-based evidence. Generative models such as MVAE (khattar2019mvae) and GSFND (tong2025generate) align multimodal features or augment training data. However, these methods often rely on domain labels or engagement signals, incur high cost, and lack interpretability, robustness, and evidence-grounded reasoning in mismatched or unseen domains.
(iii) Advance Language Model-based techniques: LLMs have advanced fake news detection through reasoning-augmented and robustness-driven strategies. GLPN-LLM (hu2025synergizingllmsgloballabel) incorporates global label reasoning, while PCoT (modzelewski2025pcotpersuasionaugmentedchainthought) augments chain-of-thought prompting with persuasion cues. Robustness has been pursued via hallucination informed reasoning (IMRRF) (li-etal-2025-imrrf), adversarial style generation (AdStyle) (10.1145/3696410.3714569), and laundering-resistant detection (das2025fakenewsdetectionllm). Retrieval and collaboration approaches include self-reflective retrieval (SR3) (Zhou_Zhang_Tan_Zhang_Li_2025) and multi-agent LLM–SLM co-learning (10704605). Data augmentation methods such as GSFND (tong2025generate) and TripleFact (xu2025triplefact) employ generative sampling and evidence-grounded triples. Yet, LLM-based systems remain limited by high computational cost, shallow multimodal integration, vulnerability to data contamination, and difficulties adapting to rapidly evolving misinformation.
Positioning of our Paper: To overcome the constraints of current approaches, we propose AMPEND-LS, an evidence-grounded, multi-persona agentic framework for robust fake news detection. Unlike domain adaptation methods that depend on domain labels or user engagement, and multimodal models limited by shallow fusion or costly encoders, AMPEND-LS explicitly integrates textual, visual, and contextual evidence with structured reasoning. A composite reliability score fuses lexical–semantic similarity, domain credibility, and temporal consistency, providing fine-grained veracity signals. Multi-persona agentic reasoning enables iterative claim validation through diverse expert roles, generating transparent justifications. This reasoning is distilled into a hybrid classification pipeline, where an LLM classifier captures rich contextual memory and a lightweight SLM provides efficient, uncertainty-aware verification. By combining evidence fusion, agentic reasoning, and scalable classification, AMPEND-LS achieves stronger factual grounding, interpretability, and adaptability than existing methods, positioning it as a deployable and trustworthy framework for combating misinformation in dynamic, multi-domain environments.
3. Problem Formulation
Given a news article $N$, represented as $N=(\mathcal{H},\mathcal{T},\mathcal{I})$, where $\mathcal{H}$ denotes the headline, $\mathcal{T}$ the article body text, and $\mathcal{I}$ an associated image. Each article has a ground-truth label $\mathcal{Y}\in\{REAL,FAKE\}.$
Challenges: Detecting fake news is challenging due to: (i) the multimodal nature of news, requiring joint reasoning across text and images, (ii) the presence of misleading or manipulated evidence, which complicates credibility assessment, and (iii) the need for explainability, where mere classification is insufficient without transparent reasoning.
Objective: The goal is to learn a prediction function $f:N\mapsto\mathcal{Y},$ that is not only accurate but also evidence-aware and explainable.
4. Methodology
Our proposed framework, AMPEND-LS (Figure 1), integrates textual, visual, and contextual cues into a unified pipeline designed to (i) handle multimodal inputs, (ii) ground predictions in credible evidence, and (iii) provide explainable reasoning.
The process begins with preprocessing, where headlines, article text, and associated images are refined into structured representations via claim extraction, text cleaning, and image summarization. An evidence retrieval and credibility assessment module then enriches the context with corroborating or refuting information from reliable sources, reducing reliance on potentially misleading inputs. At the core lies the Multi-step Multi-persona Agentic Framework (MMAF), where LLM-powered agents adopt distinct personas (e.g., journalist, supervisor, legal analyst, scientific expert) and iteratively analyze claims through Q&A interactions. This multi-perspective reasoning enables the capture of signals such as credibility, bias, and contextual consistency, producing both predictions and interpretable justification trails. Uncertain cases are not discarded but revisited with persuasion strategies to refine label quality and robustness. Finally, to ensure efficiency, an SLM adaptation stage distills LLM-generated pseudo-labels and reasoning into a lightweight model. This preserves interpretability while enabling fast, resource-efficient inference with strong generalization.
4.1. Preprocessing
The stage processes heterogeneous news inputs into structured, semantically rich representations that enable effective evidence retrieval and reasoning.
(i) Article Text Preprocessing: Raw text $\mathcal{T}$ often contains artifacts (irregular punctuation, whitespace, noise). We apply standard cleaning operations to obtain a noise-free factual sequence $\mathcal{T}_{p}$, improving claim extraction and retrieval accuracy.
(ii) Claim Generation: Since headlines $\mathcal{H}$ are concise and $\mathcal{T}_{p}$ may contain extraneous details, we distill them into a concise, entity-aware claim $\mathcal{C}=f_{1}(\mathcal{H},\mathcal{T}_{p}),$ where $f_{1}$ is an LLM-based claim generator that emphasizes salient entities and contextual cues. $\mathcal{C}$ facilitates precise evidence retrieval.
(iii) Image Summary Generation: To capture visual signals, we extract image entities $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ (e.g., objects, locations, text) using Google Cloud Vision API (https://cloud.google.com/vision). These are embedded into an LLM prompt to generate summary $\mathcal{I}_{s}=f_{2}(\mathcal{I},\mathcal{V}),$ where $f_{2}$ generates a coherent, human-readable description integrating $\mathcal{V}$ with salient visual content. $\mathcal{I}_{s}$ provides a structured textual representation of $\mathcal{I}$, suitable for fact-checking.
The preprocessing stage yields $(\mathcal{T}_{p},\mathcal{C},\mathcal{I}_{s})$, which together provide semantically rich, analyzable inputs for later stages.
4.2. Evidence and Facts Retrieval
This module retrieves and validates external information to support or refute the generated claim $\mathcal{C}$, yielding structured evidence $\mathcal{E}$ for downstream agentic reasoning.
(i) Evidence Selection: To retrieve candidate evidence, we perform a web search on the generated claim $\mathcal{C}$ using the Google Cloud Custom Search API (https://console.cloud.google.com), yielding a document set $\mathcal{D}=\{d_{i}\}_{i=1}^{n}$. To restrict the search space, we retain only the top-$K_{1}$ domains, denoted $top\text{-}K_{1}(\mathcal{D})$. For each candidate document, we compute two complementary similarity scores:
(1) $\scriptsize s_{1}=\sigma~\!\bigl(BM25(\mathcal{C},top\text{-}K_{1}(\mathcal{D}))\bigr),\qquad s_{2}=BS(\mathcal{C},top\text{-}K_{1}(\mathcal{D})),$
where, $s_{1}$ measures lexical relevance via BM25 (bm25), $s_{2}$ measures semantic similarity via BERT-Score (bert_score), and $\sigma$ is a normalization function. Formally, the evidence retrieval operator is defined as:
$f_{\text{retr}}:\mathcal{C}\;\mapsto\;\{(d_{i},s_{1}(d_{i}),s_{2}(d_{i}))\}_{i=1}^{K_{1}}.$
This hybrid design balances lexical and semantic matching, ensuring that retrieved evidence is both textually aligned and contextually relevant for claim verification.
(ii) Domain Credibility Analysis: Textual similarity scores alone cannot ensure the trustworthiness of evidence, since unreliable or biased sources may exhibit high lexical or semantic overlap with the claim. To address this, we incorporate domain credibility as an additional signal of reliability.
We adopt the Media Bias/Fact Check (MBFC) database (mediabiasfactcheck.com), which provides widely used independent assessments of media outlets by combining factual reporting quality with political/ideological bias ratings. MBFC is chosen for its comprehensive coverage and transparent methodology, making it a practical credibility reference.
Each candidate domain is assigned a categorical credibility score DCS $\in\{\textit{High},\textit{Medium},\textit{Unknown},\textit{Low}\}$, which we map to weights:
(2) $\scriptsize s_{3}(d_{i})=\{\text{High}\!\mapsto\!\alpha_{h},\;\text{Medium}\!\mapsto\!\alpha_{m},\;\text{Unknown}\!\mapsto\!\alpha_{u},\;\text{Low}\!\mapsto\!\alpha_{l}\}.$
with ordering constraints: $\alpha_{h}>\alpha_{m}>\alpha_{u}>\alpha_{l},~~\alpha_{j}\in[0,1],\;j\in\{h,m,u,l\}.$ Typical settings are $\alpha_{h}\approx 1$, $\alpha_{l}\approx 0$, and $\alpha_{u}\approx 0.5$ to avoid unfair exclusion of less popular domains. The credibility operator is therefore defined as:
$f_{\text{cred}}:top\text{-}K_{1}(\mathcal{D})\;\mapsto\;\mathbf{s}_{3}=[s_{3}(d_{1}),\dots,s_{3}(d_{K_{1}})],$
which penalizes evidence from low-credibility sources while prioritizing reputable domains for claim verification.
(iii) Temporal Evidence Filtering: This step aims to retain evidence contemporaneous with the news event while filtering out outdated reports of similar past events. In the absence of a known publication date for the target article, we infer the temporal context of the event from the retrieved evidence itself. The key idea is that relevant sources should cluster within a dominant time window, whereas outdated reports of similar past events typically fall outside this cluster. To identify this window robustly, we use the median publication date $t_{\text{median}}$ of the top-$K_{1}$ retrieved documents, as it is less sensitive to outliers than the mean. Let $t_{i}$ denote the publication date of candidate document $d_{i}$ and $t_{\text{median}}$ the median date across top-$K_{1}$ retrieved documents. The temporal deviation is defined as: $\Delta_{i}=|t_{i}-t_{\text{median}}|.$ The temporal consistency score is computed as:
(3) $\scriptsize s_{4}(d_{i})=\begin{cases}\max(\alpha_{t},1-\Delta_{i}/365),&t_{i}\neq\text{unknown}\\ 0,&\text{otherwise}\end{cases}$
where, scores close to $1$ indicate strong temporal alignment with the current event. A lower bound of $\alpha_{t}$, a hyperparameter, prevents discarding slightly delayed but relevant reports, while documents without valid timestamps are excluded. The temporal filtering operator is thus defined as
$f_{\text{temp}}:top\text{-}{K_{1}}(\mathcal{D})\;\mapsto\;\mathbf{s}_{4}=[s_{4}(d_{1}),s_{4}(d_{2}),\dots,s_{4}(d_{K_{1}})].$
The objective of this filtering is to prioritize evidence that co-occurs within the event’s active propagation window, thereby ensuring temporal relevance and excluding outdated or recycled content.
(iv) Reliability Score: For each $d_{i}\in top\text{-}K_{1}(\mathcal{D})$, the reliability score integrates all the above scores:
(4) $\scriptsize\mathcal{R}(d_{i})=\lambda_{1}s_{1}(d_{i})+\lambda_{2}s_{2}(d_{i})+\lambda_{3}s_{3}(d_{i})+\lambda_{4}s_{4}(d_{i});\qquad\sum\nolimits_{j=1}^{4}\lambda_{j}=1$
where $\lambda_{j}$ are tunable weights, for $j\in\{1,2,3,4\}$. Documents are ranked by their reliability score $\mathcal{R}(d_{i})$, and the top-$K^{\prime}_{1}$ are retained as credible evidence $\mathcal{E}_{T}$ for downstream misinformation analysis.
(v) Visual Evidence: Reverse image search is employed to trace the origin and context of images, enabling the retrieval of relevant external evidence to verify or refute potential fake news. We, therefore, perform reverse image search on the query image $\mathcal{I}$ using the Google Cloud Vision API. This yields a candidate set of documents with metadata: $\langle$ image URL, article URL, web entities, article title, article summary, date of publication $\rangle,$ from which irrelevant sources are discarded. For each candidate $d_{i}$, reliability is assessed using the same four scores as in textual evidence, lexical similarity, semantic similarity, domain credibility, and temporal consistency, resulting in an overall reliability score $\mathcal{R}(d_{i})$.
The top-$K_{2}$ candidates are ranked by $\mathcal{R}(d_{i})$, and the top-$K^{\prime}_{2}$ are retained as credible image-based evidence: $\mathcal{E}_{I}=\{d_{i}\mid d_{i}\in top\text{-}K^{\prime}_{2}(\mathcal{D}_{I})\}.$
This ensures that only temporally relevant, domain-reliable, and contextually aligned visual sources contribute to downstream misinformation analysis.
(vi) Entity–Relation Evidence: To provide structured factual grounding for claim verification, we query external knowledge bases (Wikidata, DBpedia, Google Knowledge Graph) via LLMs to extract relational facts. These are represented as knowledge graph (KG) triplets $\langle s,r,o\rangle\;\;\text{with}\;\;s,o\in\mathcal{E}_{ent},\;r\in\mathcal{R}_{KG},$ where $\mathcal{E}_{ent}=\{e_{1},e_{2},\dots,e_{m}\}$ denotes the set of entities extracted from the claim $\mathcal{C}$ (e.g., persons, organizations, locations, events), and $\mathcal{R}_{KG}=\{r_{1},r_{2},\dots,r_{p}\}$ denotes the set of factual relations supported by knowledge bases (e.g., foundedBy, memberOf, locatedIn, dateOf). For example, $\langle$ NASA,   foundedBy,   U.S. Government $\rangle$. The retrieved triplets form $K_{3}$ number of entity–relation paths: $\mathcal{E}_{KG}=\{\langle s,r,o\rangle_{k}\}_{k=1}^{K_{3}}.$
These structured relations are integrated with prior evidence sets to strengthen factual validation, enabling the framework to support or refute $\mathcal{C}$ against authoritative knowledge sources.
Collectively, this module constructs the structured evidence set $\mathcal{E}=\{\mathcal{E}_{T},\mathcal{E}_{I},\mathcal{E}_{KG}\},$ This unified evidence representation $\mathcal{E}$ is forwarded to the next stage for advanced misinformation analysis.
4.3. Multi-step Multi-persona Agentic Framework (MMAF)
Building on the structured evidence $\mathcal{E}$, we introduce an MMAF for fake news detection. MMAF organizes LLM agents into distinct roles: Supervisor, Journalist, Legal Analyst, and Scientific Expert, supported by an Answering Agent. Unlike single-agent prompting, these personas iteratively pose probing questions, with the Answering Agent delivering concise, evidence-grounded responses. A shared memory accumulates these interactions, enabling coordinated reasoning. This explicit role separation, disciplined answering, and iterative enrichment yield more robust, transparent, and evidence-aware verification than conventional approaches.
Definition 4.1 (MMAF).
An MMAF for evidence-aware fake news detection is a structured system of LLM-based agents. Formally, it is defined as: $\mathcal{F}=\{A_{\text{Supervisor}},A_{\text{Journalist}},A_{\text{Legal}},A_{\text{Scientific}},A_{\text{Ans}}\}$, where each role-specific agent has the tuple structure $A_{i}=(\mathcal{O},\mathcal{A},\pi_{i}^{q}$, $\mathcal{M},\mu$, $\mathcal{G}_{i}),$ for $i\in\{\text{Supervisor},\ \text{Journalist},\ \text{Legal},\ \text{Scientific}\}.$ A separate answering agent $A_{\text{Ans}}=(\mathcal{O},\mathcal{A}_{\text{Ans}},\pi^{a},\mathcal{M},\mu,\mathcal{G}_{\text{Ans}})$.
Observation space: $\mathcal{O}=(\mathcal{N}_{e},\mathcal{K}_{\text{LLM}}),\quad\mathcal{N}_{e}=\{\mathcal{H},\mathcal{T}_{p},\mathcal{C},\mathcal{I}_{s},\mathcal{E}\},$ where, $\mathcal{H}$ = headline, $\mathcal{T}_{p}$ = preprocessed article text, $\mathcal{C}$ = extracted claim, $\mathcal{I}_{s}$ = image summary from image entities, $\mathcal{E}$ = retrieved evidence (credibility checks, reverse image search, KG facts), and $\mathcal{K}_{\text{LLM}}$ = LLM’s internal knowledge.
Action space: We distinguish between different types of actions: $\mathcal{Q}:\text{space of probing questions},~\mathcal{A}_{\text{Ans}}:\text{space of concise answers}.$. Thus, the global action space is $\mathcal{A}=\mathcal{Q}\cup\mathcal{A}_{\text{Ans}}.$
Policies: Each persona $A_{i}$ is guided by a questioning policy $\pi_{i}^{q}:\mathcal{O}\times\mathcal{M}\to\mathcal{Q},$ which maps observations and memory to a probing question $q_{i}^{t}$. Agent $A_{\text{Ans}}$ uses an answering policy $\pi^{a}:\mathcal{Q}\times\mathcal{O}\times\mathcal{M}\to\mathcal{A}_{\text{Ans}},$ which generates concise, evidence-grounded answers $a_{i}^{t}$.
Memory $\mathcal{M}$ is a structured contextual memory containing Q&A history, abstracted insights, and summaries.
Memory update function $\mu:\mathcal{M}\times(\mathcal{Q},\mathcal{A}_{\text{Ans}})\to\mathcal{M}$ updates memory by synthesizing each new Q&A pair.
Goals: Each agent has a role-specific goal:
  • $\mathcal{G}_{\text{Supervisor}}$: determine overall claim veracity,
  • $\mathcal{G}_{\text{Journalist}}$: probe credibility, bias, and contextual consistency,
  • $\mathcal{G}_{\text{Legal}}$: check regulatory or legal compliance,
  • $\mathcal{G}_{\text{Scientific}}$: validate factual and scientific correctness,
  • $\mathcal{G}_{\text{Ans}}$: generate concise, evidence-based answers to persona queries.
Iterative reasoning: The Q&A cycle iterates over personas, incrementally enriching $\mathcal{M}$. At each round $t=1,\dots,\tau$, with observation $o^{t}\sim\mathcal{O}$:
(5) $\scriptsize\begin{split}q_{i}^{t}=\pi_{i}^{q}(o^{t},M^{t-1})\in\mathcal{Q}~;\hskip 16.38895pta_{i}^{t}=\pi^{a}(q_{i}^{t},o^{t},M^{t-1})\in\mathcal{A}_{\text{Ans}}~;\\ M^{t}=\mu(M^{t-1},(q_{i}^{t},a_{i}^{t})).\end{split}$
Beyond supporting classification, the shared memory $\mathcal{M}$ regulates the Q&A process by preventing redundant questions, mitigating hallucinations, and enforcing insight diversity across personas. This ensures that MMAF systematically covers complementary verification dimensions, including factual validity, source credibility, political/regional bias, and legal compliance. The pseudo-code for the MMAF reasoning process is provided in Algorithm 1.
Algorithm 1 Multi-step Multi-persona Agentic Reasoning
1:Number of rounds $\tau$; persona set $\mathcal{P}=\{\text{Supervisor},\text{Journalist},\text{Legal},\text{Scientific}\}$;   Observation components $\mathcal{N}_{e}=\{\mathcal{H},\mathcal{T}_{p},\mathcal{C},\mathcal{I}_{s},\mathcal{E}\}$ and internal knowledge $\mathcal{K}_{\text{LLM}}$
2:Structured contextual memory $\mathcal{M}^{\tau}$
3:Define observation space $\mathcal{O}\leftarrow(\mathcal{N}_{e},\mathcal{K}_{\text{LLM}})$
4:Initialize memory $\mathcal{M}^{0}\leftarrow\textsc{Init}(\mathcal{N}_{e})$
5:Given persona question policies $\{\pi_{i}^{q}:\mathcal{O}\times\mathcal{M}\!\to\!\mathcal{Q}\}_{i\in\mathcal{P}}$ and answering policy $\pi^{a}:\mathcal{Q}\times\mathcal{O}\times\mathcal{M}\!\to\!\mathcal{A}_{\text{Ans}}$
6:for $t=1$ to $\tau$ do
7:  Sample/compose observation $o^{t}\sim\mathcal{O}$
8:  for each $i\in\mathcal{P}$ do
9:    Question: $q_{i}^{t}\leftarrow\pi_{i}^{q}(o^{t},\mathcal{M}^{t-1})$
10:    Answer: $a_{i}^{t}\leftarrow\pi^{a}(q_{i}^{t},\,o^{t},\,\mathcal{M}^{t-1})$
11:    Update memory: $\mathcal{M}^{t}\leftarrow\mu(\mathcal{M}^{t-1},\,(q_{i}^{t},a_{i}^{t}))$
12:  end for
13:end for
14:return $\mathcal{M}^{\tau}$
4.4. LLM Pseudo-Labeling: Policy, Decision Rules, and Output
We formulate pseudo-labeling as a disciplined, evidence-aware decision stage where an LLM, constrained by a policy, operates over the structured final memory built from textual, visual, and KG evidence. The objective is to produce REAL/ FAKE/ UNCERTAIN labels with calibrated confidence and a concise, auditable justification. To mitigate hallucinations and source blindness, the policy requires using the entire memory as the primary factual base, consulting general world knowledge only when necessary, and invoking optional tools (e.g., persuasion-based refinement) only under uncertainty. Evidence items are tagged as supporting or contradicting the claim and weighted by credibility and temporal relevance; the LLM aggregates these signals to choose a verdict and articulate a brief rationale that cites the memory. The output is a strict JSON record (verdict, confidence, justification), enabling downstream automation and supervision of a compact SLM while preserving transparency.
We obtain evidence-grounded pseudo-labels by querying an LLM classifier under a constrained policy:
(6) $\scriptsize(\mathcal{Y}_{p},\,\mathcal{C}_{s},\,\mathcal{J})\;=\;\Phi~\!\bigl(~\pi_{c};\;\mathcal{M}^{\tau},\,{\text{FewShots}}~\bigr),$
where, $\mathcal{Y}_{p}\in\{\textit{REAL},\textit{FAKE},\textit{UNCERTAIN}\}$, $\mathcal{C}_{s}\in[0,1]$ is a confidence score, and $\mathcal{J}$ is a concise justification that cites items in $\mathcal{M}^{\tau}$ (primary source) for interoperability. FewShots guidance improves decision consistency; memory citation ensures auditability.
To handle cases where aggregated evidence is inconclusive, we apply a persuasion-based refinement module only to UNCERTAIN instances. We first detail the persuasion-based refinement, then return to the LLM policy decision rules.
4.4.1. Persuasion-based Refinement
We formalize persuasion-based refinement as a gated, evidence-augmenting procedure applied only to items initially labeled uncertain by the LLM pseudo-labeler.
Gating: Let the uncertain set be defined as:
(7) $\scriptsize\mathcal{U}=\{N\mid\mathcal{Y}_{p}(N)=\textit{UNCERTAIN}=\Phi~\!\bigl(~\pi_{c};\;\mathcal{M}^{\tau},\,\text{FewShots}~\bigr)\}.$
This module is activated only when the LLM classifier cannot reach a stable verdict, minimizing unnecessary procedure. We begin by defining the Persuasion Knowledge Model (PKM) agent.
Definition 4.2 (PKM Agent).
A PKM agent is an auxiliary LLM-based agent that detects and summarizes persuasion techniques to refine uncertain pseudo-labels. It is defined as: $A_{\text{PKM}}=\bigl(\mathcal{O},\,\mathcal{A}_{\text{PKM}}$, $\pi_{\text{PKM}},\,\mathcal{M},\,\mu,\,\mathcal{G}_{\text{PKM}}\bigr).$
Observation space: $\mathcal{O}=(\mathcal{N}_{e},\,\mathcal{K}_{\text{LLM}}),~~\mathcal{N}_{e}=\{\mathcal{H},\mathcal{T}_{p},\mathcal{C},\mathcal{I}_{s},\mathcal{E}\}.$ PKM primarily consumes $\mathcal{N}_{e}^{\text{PKM}}=\{\mathcal{C},\mathcal{T}_{p}\}$ together with the current memory ${\mathcal{M}}^{\tau}$ (which may already encode textual, visual, KG evidence and MMAF knowledge).
Action space: $\mathcal{A}_{\text{PKM}}=\Bigl\{\,z\in\{0,1\}^{T},~\ \mathcal{S}=\{s_{t}\},~\ u\in\mathbb{N}^{G},~\ \Pi\in[0,1],~\ \mathcal{J}_{\text{PKM}}\Bigr\}$. Here, $T=23$ persuasion techniques (JRC taxonomy (piskorski2023news)) and $G=6$ categories (justification, simplification, distraction, call, reputation attack, and manipulative wording); $z$ is the technique-activation vector, $\mathcal{S}$ is extracted evidence spans (phrases), $u$ is category counts, $\Pi$ is a normalized persuasion index, and $\mathcal{J}_{\text{PKM}}$ is a concise narrative summary.
Policy: $\pi_{\text{PKM}}:\ \mathcal{O}\times\mathcal{M}\ \longrightarrow\ \mathcal{A}_{\text{PKM}}$. Given $(\mathcal{C},\mathcal{T}_{p},\mathcal{I}_{s},\mathcal{M})$, the policy produces (i) $z,\mathcal{S}$ via detection over content; (ii) $u=\mathscr{A}z$ using a fixed category–technique incidence matrix ${\mathscr{A}}\in\{0,1\}^{G\times T}$; (iii) $\Pi=\frac{\beta^{\top}u}{\|\beta\|_{1}}$ with $\beta\in\mathbb{R}_{+}^{G}$ denoting the importance of the category; (iv) $\mathcal{J}_{\text{PKM}}=\mathrm{Summarize}(z,\mathcal{S})$. The PKM agent is gated, it is invoked only if ${g_{\text{PKM}}}={\mathbb{1}}[\mathcal{Y}_{p}=\textit{UNCERTAIN}]$; $\mathbb{1}$}[.] is the indicator function.
Memory: $\mathcal{M}$ stores structured Q&A history, abstracted insights, and evidence summaries.
Memory update: The update synthesizes persuasion signals into memory for downstream classification: $\mathcal{M}^{+}=\mu~\!\bigl(\mathcal{M},\mathcal{A}_{\text{PKM}}\bigr).$
Goal: To detect and explain persuasion to refine uncertain decisions and improve interoperability $\mathcal{G}_{\text{PKM}}$.
Activation and effect: When $g_{\text{PKM}}=1$, PKM runs once (i.e., no Q&A loop), updates $\mathcal{M}\to\mathcal{M}^{+}$, after which persuasion-aware re-scoring or re-querying of the pseudo-labeler is performed.
4.4.2. LLM Pseudo-Labeling Policy
We formalize pseudo-labeling as a constrained decision function operating on the final memory. Let $\mathcal{M}^{\tau}$ be the structured memory, and let $\mathcal{E}^{+}$$\mathcal{E}^{-}$ denotes supporting/contradicting evidence items, each endowed with a reliability score ${\mathcal{R}}(d_{i})$ for $d_{i}\in{\mathcal{E}}_{T}\cup{\mathcal{E}}_{I}$.
Policy constraints: We define a two-stage decision with strict source and tool gating:
${\Phi}:\ \mathcal{M}^{\tau}\ \mapsto\ ({\mathcal{Y}}_{p},{\mathcal{C}}_{s},{\mathcal{J}})~~;\quad\tilde{\Phi}:\ \mathcal{M}^{+}\mapsto\ (\mathcal{Y}_{p},\mathcal{C}_{s},\mathcal{J}),\text{ if }g_{\text{PKM}}=1$
where, the preliminary decision $\Phi$} must use only $\mathcal{M}^{\tau}$. The refined decision $\tilde{\Phi}$ may additionally consult persuasion-based refinement, for uncertain cases. The rationale is that hard gating reduces hallucinations and tool-induced drift by prioritizing curated memory as the primary knowledge source, while auxiliary tools are activated only when principled uncertainty arises.
Evidence-weighted aggregation and confidence: We aggregate reliability-weighted support/contradiction as:
(8) $\scriptsize S^{+}=\sum_{d_{i}\in\mathcal{E}^{+}}\mathcal{R}(d_{i});\qquad S^{-}=\sum_{d_{i}\in\mathcal{E}^{-}}\mathcal{R}(d_{i});\qquad$
Weighting by source reliability yields robust signal integration.
Preliminary decision rule $\tilde{\mathcal{Y}}_{p}$: Let $\mathcal{E}^{\pm}_{H}$ be the subsets of high-reliable evidence (i.e., reliability score is more than a threshold $\alpha_{h}$) within $\mathcal{E}^{\pm}$; let $\gamma,\eta>0$ be hyperparameters.
(9) $\scriptsize\tilde{\mathcal{Y}}_{p}\;=\;\begin{cases}\textit{REAL},&\left(|\mathcal{E}^{+}_{H}|>0\wedge|\mathcal{E}^{-}_{H}|=0\right)\ \ \text{or}\\[1.0pt] &\left(|\mathcal{E}^{+}_{H}|+|\mathcal{E}^{-}_{H}|=0\wedge S^{+}_{\text{Med}}-S^{-}_{\text{Med}}\geq\gamma\right),\\[4.0pt] \textit{FAKE},&\left(|\mathcal{E}^{-}_{H}|>0\wedge|\mathcal{E}^{+}_{H}|=0\right)\ \ \text{or}\\[1.0pt] &\left(\text{only low-reliable support}\right)\ \ \text{or}\ \ \left(\psi_{\text{impl}}(N)=1\right),\\[4.0pt] \textit{UNCERTAIN},&\left(|\mathcal{E}^{+}_{H}|>0\wedge|\mathcal{E}^{-}_{H}|>0\right)\ \ \text{or}\ \ \left(S^{+}+S^{-}<\eta\right),\end{cases}$
where, $S^{\pm}_{\text{Med}}$ denote the medium-reliable components (i.e., reliability score is between $\alpha_{h}$ and $\alpha_{m}$) of $S^{\pm}$, and $\psi_{\text{impl}}(N)\in\{0,1\}$ flags implausible/sensational content by deterministic checks. High-reliable precedence encodes epistemic dominance; medium-only consistency covers cases with no high-reliable sources; uncertainty guards against conflicted or sparse evidence.
The module outputs a structured triplet in the form (verdict, confidence, justification); $\texttt{verdict}:\mathrm{map}(\mathcal{Y}_{p})\in\{\textit{REAL},\textit{FAKE}$, $\textit{UNCERTAIN}\}$, $\texttt{confidence}=\mathrm{round}\!\left(100\,\mathcal{C}_{s}\right)$, $\texttt{justification}=\mathcal{J}$, with $\mathcal{J}$ restricted to cite items from $\mathcal{M}^{\tau}$ (and explicitly mark auxiliary knowledge, if any). Deterministic rendering ensures auditability and supports SLM training.
4.5. SLM Classifier for Final Output
The LLM-based pseudo-labeling stage provides evidence-grounded reasoning and interpretability, but deploying LLMs alone is prone to overfitting on superficial linguistic cues. To balance interpretability with efficiency, we employ an SLM classifier that consumes both raw news content and the reasoning artifacts generated by the LLM, thus integrating factual content with structured justifications. Formally, let the structured input sequence be:
(10) $\scriptsize x=\big[\mathcal{H}\;\oplus\;\langle\texttt{SEP}\rangle\;\mathcal{T}_{p}\;\oplus\;\langle\texttt{SEP}\rangle\;\mathcal{I}_{s}\;\oplus\;\langle\texttt{SEP}\rangle\;\mathcal{J}\big],$
where, $\mathcal{H}$ is the headline, $\mathcal{T}_{p}$ is preprocessed article text, $\mathcal{I}_{s}$ is the image summary, $\mathcal{J}$ is the justification from the LLM classifier, and $\oplus$ denotes concatenation with structural separators $\langle\texttt{SEP}\rangle$. This ensures modality-aware encoding and preserves logical boundaries between different information sources.
We instantiate the classifier as a pre-trained transformer encoder $f_{\theta}$ (DistilRoBERTa-base) parameterized by $\theta$. The classifier outputs the final decision label: $\mathcal{Y}_{s}=f_{\theta}(x)\in\{\textit{REAL},\textit{FAKE}\}.$
The parameters $\theta$ are optimized using standard cross-entropy loss. This design ensures (i) efficiency, since $f_{\theta}$ is lightweight compared to LLMs, (ii) robustness, as the SLM benefits from the distilled reasoning $\mathcal{J}$ rather than overfitting to raw text alone, and (iii) generalization, as the combined representation $(\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s},\mathcal{J})$ anchors decisions in both factual content and structured evidence.
4.6. Model Deployment
In inference, each news item undergoes preprocessing to yield structured content and evidence. These inputs are evaluated by a multi-persona reasoning module, which builds a consolidated memory passed to an LLM classifier. The LLM outputs preliminary labels with justifications, ensuring interpretability. For uncertain cases, a persuasion-based refinement module identifies manipulative cues to reduce ambiguity. Finally, the enriched representation, combining raw content and reasoning outputs, is processed by an SLM classifier, which produces the final decision. This modular design balances evidence-grounded reasoning (via LLM), disambiguation (via persuasion refinement), and computational efficiency with generalization (via SLM).
5. Experiments
All experiments were conducted on a system equipped with PyTorch 2.5.1, CUDA 12.1, NVIDIA A100-SXM4-40GB GPU, AMD EPYC 7742 64-core CPU with 1 TB RAM. We used the Google Cloud Custom Search API to retrieve external evidence, and the OpenAI GPT-4o mini for LLM-based inference. For the final verdict, DistilRoBERTa-base (distilroberta-base) was finetuned as SLM to obtain REAL or FAKE.
5.1. Datasets and Evaluation Metrics
Datasets: To evaluate our model, we used three widely adopted fake news datasets: PolitiFact (shu2020fakenewsnet), GossipCop (shu2020fakenewsnet), and MMCoVaR (chen2021mmcovar). PolitiFact contains fact-checked political claims with veracity labels, while GossipCop consists of celebrity and entertainment-related news articles, annotated as REAL or FAKE. MMCoVaR is a multimodal dataset on COVID-19 vaccine-related misinformation, comprising both textual posts and associated images. The detailed statistics of all datasets are presented in Table 1. Brief statistics of these datasets are presented in Table 1. For training and evaluation, we used the train-validation-test splits in a ratio of 7:2:17{:}2{:}1.
Table 1. Statistics of the datasets
Statistics PolitiFact GossipCop MMCoVaR
# Fake News 103 250 706
# Real News 172 250 1530
# Images 275 500 2236
Evaluation Metrics: To evaluate the effectiveness of the model, we employ four widely adopted metrics: Accuracy % ($Acc$), F1 Score % ($F1$), Precision % ($Pre$), and Recall % ($Rec$). In addition, we report an improvement metric that measures the percentage gain in our method over the second-best baseline. To evaluate the quality of the generated rationale (LLM justification $\mathcal{J}$), we employ three complementary metrics: (i) the Utility metric $\mathcal{U}$, capturing the effectiveness of the rationale as the performance difference between fine-tuned SLMs trained with and without LLM justification $\mathcal{J}$; (ii) the Leakage-Adjusted Simulatability ($LAS$) metric (las_metric), assessing faithfulness, how well $\mathcal{J}$ aids in simulating the verdict without leaking label information, computed as $\mathbb{1}[\mathcal{Y}_{s}\mid\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s},\mathcal{J}]-\mathbb{1}[\mathcal{Y}_{s}\mid\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}];$ where $\mathbb{1}[\cdot]$ is an indicator function that equals $1$ if the model correctly predicts the true label and 0 otherwise; and (iii) the Rationale Evaluation via conditional $\mathcal{V}$-information ($REV$) metric (rev_metric), quantifying informativeness, the additional label-relevant information provided by $\mathcal{J}$, computed as the log-likelihood difference of the true label $\mathcal{Y}$ between two models $f_{\theta}$ and $f^{\prime}_{\theta}$ trained with $\mathcal{J}$ and without $\mathcal{J}$ as inputs, respectively: $\log f_{\theta}([\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}],\mathcal{J})(\mathcal{Y})-\log f^{\prime}_{\theta}([\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}])(\mathcal{Y}).$ Each metric is averaged over all samples in the test dataset.
5.2. Results and Analysis
Table 2. Comparison with SOTA methods
Dataset Method AccAcc $F1$ PrePre RecRec PolitiFact LMDD (Kheddache2025) 80.86 75.99 80.05 72.34 FactAgent (Li2024) 81.41 81.09 80.79 81.41 MFCL (CHEN2025112800) 81.82 82.52 81.61 83.47 BREAK (yin2025break) 82.14 82.03 81.79 82.29 MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) 82.14 81.10 82.14 80.09 EARAM (10.1145/3696410.3714532) 82.93 81.11 80.90 81.32 FakingRecipe (10.1145/3664647.3680663) 83.33 80.07 80.83 79.44 AMPEND-LS (Proposed) 92.18 ± 1.93 91.64 ± 2.12 92.06 ± 1.85 91.40 ± 2.43 Improvement (%) 10.62 11.05 12.08 9.50 GossipCop LMDD (Kheddache2025) 65.17 71.05 78.21 65.10 FactAgent (Li2024) 78.04 77.95 78.45 77.50 FakingRecipe (10.1145/3664647.3680663) 80.26 80.27 80.21 80.26 MFCL (CHEN2025112800) 81.81 81.35 81.65 81.06 BREAK (yin2025break) 82.00 82.11 82.51 81.73 EARAM (10.1145/3696410.3714532) 85.00 85.49 86.00 85.00 MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) 88.00 84.09 88.31 80.26 AMPEND-LS (Proposed) 93.43 ± 1.53 93.24 ± 1.59 93.44 ± 1.48 93.29 ± 1.55 Improvement (%) 6.17 9.07 5.81 9.75 MMCoVaR LMDD (Kheddache2025) 64.62 63.54 78.61 53.40 FactAgent (Li2024) 79.14 73.70 74.45 73.10 FakingRecipe (10.1145/3664647.3680663) 79.96 78.61 79.35 77.92 MFCL (CHEN2025112800) 82.06 78.19 77.82 78.57 MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) 90.00 88.34 88.19 88.48 EARAM (10.1145/3696410.3714532) 90.36 88.51 88.32 88.70 BREAK (yin2025break) 91.89 89.27 88.73 89.83 AMPEND-LS (Proposed) 96.46 ± 1.11 95.56 ± 1.31 96.30 ± 1.48 94.96 ± 1.55 Improvement (%) 4.97 7.05 8.53 5.71
To benchmark our approach, AMPEND-LS, we compare it against two sets of baselines, (i) LLM-based methods: LMDD (Kheddache2025) and FactAgent (Li2024), and (ii) Learning-based methods, including five recent state-of-the-art (SOTA) methods, FakingRecipe (10.1145/3664647.3680663), MFCL (CHEN2025112800), MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025), BREAK (yin2025break), and EARAM (10.1145/3696410.3714532).
Table 2 shows that AMPEND-LS consistently outperformed the SOTA baselines in all three datasets: PolitiFact, GossipCop, and MMCoVaR, achieving the highest accuracy, $F1$, precision, and recall simultaneously. The largest improvements were observed on PolitiFact (+10.62% $Acc$, +11.84% $F1$), highlighting its ability to handle politically nuanced misinformation, while steady improvements on GossipCop and MMCoVaR confirm strong generalization to both text-heavy and multimodal inputs. Unlike baselines that trade off precision (e.g., MGCA) or recall (e.g., MFCL), AMPEND-LS achieved balanced improvements across all metrics, attributable to its complementary modules of agentic reasoning, persuasion-based refinement, and credibility-guided evidence fusion. Competing methods, such as BREAK, performed well only in specific domains (e.g., MMCoVaR) but failed to generalize. The tight alignment of accuracy, precision, recall, and $F1$ in AMPEND-LS further indicates stable decision boundaries and reduced overfitting. Collectively, these results validate AMPEND-LS as a robust, adaptable, and evidence-grounded framework capable of handling diverse misinformation domains with transparency and resilience.
5.3. Ablation Study
To assess the contribution of individual components in AMPEND-LS, we conduct a systematic ablation on PolitiFact and GossipCop, with results shown in Table 3. The key findings are:
Table 3. Module-wise ablation study
Dataset Method AccAcc $F1$ PrePre RecRec
PolitiFact PKM 73.13 73.49 81.23 67.10
Evidence 74.26 73.16 79.12 68.05
Agentic 79.92 78.16 81.40 75.17
w/o PKM 85.77 86.28 87.31 85.28
ASPEND 84.13 84.18 87.32 81.26
AMPEND 88.54 88.24 90.26 86.32
AMPEND-LS 92.18 91.64 92.06 91.40
GossipCop PKM 67.90 66.67 72.02 62.07
Evidence 79.67 79.71 80.06 79.37
Agentic 83.14 83.65 84.20 83.11
w/o PKM 86.28 86.03 86.13 85.98
ASPEND 88.80 89.01 89.08 88.92
AMPEND 92.68 92.56 93.04 92.10
AMPEND-LS 93.43 93.24 93.44 93.29
(i) Impact of Persuasion Analysis (PKM): PKM alone achieved relatively weak performance ($Acc$: 73.13 on PolitiFact, 67.90 on GossipCop), confirming that persuasion cues by themselves cannot reliably classify misinformation. However, ablating PKM from the full system (i.e., w/o PKM) causes notable drops (e.g., –20.66% $Acc$ on PolitiFact, –37.60% on GossipCop), highlighting its value in resolving uncertain cases where factual evidence is ambiguous. Thus, PKM is complementary: it does not excel standalone but adds critical discriminative power when integrated.
(ii) Impact of Evidence Base: Evidence-only models ($Acc$: 74.26 on PolitiFact, 79.67 on GossipCop) establish factual signals as a solid baseline but fall short compared to more integrated approaches. The large gains when evidence was combined within AMPEND (e.g., +20.61% $F1$ on PolitiFact) show that credibility scoring, temporal filtering, and multimodal grounding are indispensable for anchoring agentic reasoning to verifiable facts.
(iii) Impact of Agentic Reasoning: The standalone agentic framework outperformed PKM and Evidence individually ($Acc$ 79.92 PolitiFact, 83.14 GossipCop), demonstrating its ability to structure reasoning and probe claims. However, without external modules, it still underperformed compared to the full system. This indicates that while reasoning agents can organize analysis, they need persuasion and evidence signals for robustness.
(iv) ASPEND vs. AMPEND: A single-persona reasoning model, ASPEND, achieved strong performance ($Acc$ 84.13 PolitiFact, 88.80 GossipCop), but lagged behind the multi-persona AMPEND ($Acc$: 88.54 PolitiFact, 92.68 GossipCop). This confirms that multi-persona reasoning captures diverse perspectives (legal, journalistic, scientific, supervisory) that a single persona cannot, leading to more reliable and nuanced verdicts.
(v) AMPEND vs. AMPEND-LS: The final hybrid system AMPEND-LS, integrating the complementary SLM, yields further improvements ($Acc$: 92.18 PolitiFact, 93.43 GossipCop). The gains, though modest over AMPEND, reflect significant stability: the SLM acts as a lightweight decision layer that mitigates LLM uncertainty and hallucinations, ensuring balanced precision–recall and reducing overfitting. This synergy between LLM reasoning and SLM classification is crucial for both scalability and reliability.
12348585878789899191RoundsPerformance (%)AccAccF1F1PrePreRecRec
Figure 2. Performance across multi-rounds Q&A on PoltiFact.
5.4. Sensitivity of the Number of Rounds for Efficient Multi-Persona Reasoning
Figure 2 shows that AMPEND-LS exhibits stable yet incremental performance gains across multiple rounds of multi-persona Q&A on PolitiFact: $Acc$ improves from 88.54 at round-1 to 89.30 at round-4 (+0.76%), $F1$ from 87.34 to 88.54 (+1.20%), $Rec$ from 86.32 to 87.82 (+1.50%), while $Pre$ remains consistently high ($\sim$90%) with minimal fluctuations. These results highlight that each additional round contributes positively by refining evidence and balancing precision–recall trade-offs; however, the marginal improvements diminish after round 3. Since multi-agent Q&A cycles rely on costly API calls to LLMs, continuing indefinitely is neither practical nor efficient. Thus, a sweet spot can be identified around 2–3 rounds, where significant gains in accuracy and $F1$ are already realized, while avoiding unnecessary computational and monetary overhead. This balance between performance improvements and resource efficiency underscores the practicality of limiting reasoning depth while preserving robustness in real-world deployments.
5.5. Information-Theoretic Evaluation of LLM-Justification
To assess the effectiveness of LLM-generated justifications, we evaluate their impact across three complementary dimensions: $REV$ (rev_metric), $LAS$ (las_metric), and the utility performance metrics $\mathcal{U}_{i}$, where $i\in\{\text{$Acc$},\text{$F1$},\text{$Pre$},\text{$Rec$}\}$, as reported in Table 4. Across all datasets, our model exhibits consistently positive $REV$ and $LAS$ scores, demonstrating that LLM justifications contribute additional semantic information (higher $REV$) while maintaining faithful alignment with model predictions (positive $LAS$). More importantly, the utility metrics show notable performance gains, indicating that integrating LLM rationales into the SLM finetuning effectively enhances both predictive accuracy and explanatory reliability.
Among all datasets, PolitiFact achieves the largest improvements, suggesting that factual contexts benefit more from explicit justification signals, potentially due to the causal alignment between evidence and claims. GossipCop and MMCoVaR also show steady improvements, highlighting the AMPEND-LS adaptability across different content domains and rationale granularities. These results underscore that faithful and informative LLM justifications can meaningfully improve downstream task performance, bridging the gap between explainability and utility in aligned SLM models.
Table 4. LLM justification analysis
Dataset REVREV LASLAS $\mathcal{U}_{\text{$Acc$}}$ $\mathcal{U}_{\text{$F1$}}$ $\mathcal{U}_{\text{$Pre$}}$ $\mathcal{U}_{\text{$Rec$}}$ PolitiFact 0.3885 0.0976 9.76 11.72 6.58 13.62 GossipCop 0.2460 0.0789 7.89 7.98 6.95 7.89 MMCoVaR 0.1845 0.0363 3.63 4.90 3.09 6.04
6. Limitations and Future Work
While AMPEND-LS achieves strong performance and interpretability across diverse benchmarks, several limitations highlight opportunities for future extensions. The framework’s reliance on external APIs (e.g., web and reverse image search) may introduce latency, coverage gaps, or dependency risks; future work can investigate retrieval-agnostic or locally cached evidence pipelines to ensure robustness. The multi-persona reasoning mechanism, though improving accuracy and transparency, requires repeated LLM interactions that increase computational and monetary cost. Designing adaptive cost-aware policies to dynamically control reasoning depth offers a promising direction for scalable real-time deployment. Moreover, our experiments are limited to three English-centric datasets (PolitiFact, GossipCop, MMCoVaR). Extending evaluation to multilingual, low-resource, and culturally diverse misinformation scenarios is a critical step toward broader applicability. Current credibility scoring draws on third-party databases such as MBFC, which may not capture evolving global media ecosystems; adaptive credibility estimation that updates with new outlets can strengthen long-term resilience. Finally, while persuasion detection and evidence fusion enhance robustness, adversarial manipulations such as AI-generated deepfakes or coordinated campaigns remain an open frontier. Incorporating adversarial training, generative counterfactuals, and stronger multimodal alignment will further improve resilience.
By acknowledging these challenges as natural extensions, we emphasize that AMPEND-LS provides a solid and adaptable foundation. Addressing the above directions will not only enhance scalability and inclusivity but also strengthen the system’s role as a deployable, trustworthy solution for combating misinformation in dynamic, real-world environments.
7. Conclusion
This work introduced AMPEND-LS, an agentic multi-persona framework for evidence-aware fake news detection that unifies multimodal evidence retrieval, structured reasoning, and hybrid classification. By combining LLM-driven multi-perspective analysis with lightweight SLM adaptation, the framework achieves state-of-the-art performance while offering transparent and scalable decision-making. Extensive experiments across political, entertainment, and health misinformation domains confirm its robustness and interpretability. Looking ahead, the modular design provides a strong foundation for advancing toward multilingual, low-resource, and adversarially robust misinformation defense, thereby paving the way for trustworthy deployment in real-world, high-stakes information ecosystems.
Appendix A Configuration Details
Table 5 presents the hyperparameters of AMPEND-LS. The top document counts ($K_{i}$ or $K^{\prime}_{i}$) and the number of KG paths are set empirically to maintain relevance while reducing noise. Domain credibility thresholds align with MBFC category labels, and the temporal filter $\alpha_{t}$ excludes documents published within roughly a year. The number of persona Q&A steps is tuned using the PolitiFact dataset, while the LLM temperature is adjusted to balance creative generation with focused inference.
Table 5. Model configuration and hyperparameters
Attribute Value Top document Counts $K_{1}=K_{2}=15,K^{\prime}_{1}=K^{\prime}_{2}=10$ Domain Credibility Score $\alpha_{h}=0.9,\alpha_{m}=0.7,\alpha_{u}=0.5,\alpha_{l}=0.3$ Temporal Evidence Filtering $\alpha_{t}=0.2$ # of KG Triplet 10 # of Rounds ($\tau$) 4 Temperature Claim generation 0.30 KG generation 0.20 Persuasion generation 0.30 Question generation 0.70 Answer generation 0.40 Memory Update 0.50 LLM Classifier 0.25
Table 6. Qualitative Analysis: Correctly Classified Cases
Image Headline Article Content LLM Justification Ground Truth [Uncaptioned image] Air India Flight 171 Air India Flight 171 was a scheduled passenger flight from Ahmedabad Airport in India to London Gatwick Airport in the United Kingdom that crashed 32 seconds after takeoff at 13:39 IST (08:09 UTC) on 12 June 2025. All 12 crew members and 229 of the 230 passengers aboard died. On the ground, 19 people were killed and 67 others were seriously injured…. Verdict: REAL; Confidence: 95 Justification: The claim regarding the crash of Air India Flight 171 on June 12, 2025, is confirmed by multiple credible sources, including a preliminary report from India’s Aircraft Accident Investigation Bureau (AAIB) and coverage from medium-credibility sources like Wikipedia (wikipedia.org) and Aerotime (aerotime.aero) . These sources provide specific details about the crash, including the loss of thrust due to fuel control switches moving to CUTOFF, corroborating the tragic loss of life and the ongoing investigation. REAL [Uncaptioned image] Cristiano Ronaldo Joining Lionel Messi at Inter Miami for Club World Cup? In a surprising turn of events, football legends Cristiano Ronaldo and Lionel Messi are reportedly set to team up at Inter Miami for the upcoming Club World Cup. Sources close to the club have hinted at a groundbreaking partnership between the two superstars, aiming to elevate the team’s performance on the global stage… Verdict: FAKE; Confidence: 95
Justification: Cristiano Ronaldo remains under contract with Al Nassr, and there is no credible evidence supporting his transfer to Inter Miami. Reports suggesting this potential move are speculative and lack official confirmation from reputable sources. Messi’s confirmed presence at Inter Miami is the only verified aspect of the claim. Key sources include: (ESPN) and (NPR).
FAKE
Table 7. Qualitative Analysis: Misclassified Cases
Image Headline Article Content LLM Justification Ground Truth [Uncaptioned image] Ohio student suspended for staying in class during National Walkout Day An Ohio high school student said he tried to stay apolitical during the National Walkout Day over gun violence but was suspended for his choice to remain in a classroom instead of joining protests or the alternative, going to study hall… Verdict: REAL; Confidence: 95 Justification: Multiple high-credibility sources, including Snopes (snopes.com) and Politifact (politifact.com), confirm that Jacob Shoemaker was indeed suspended for remaining in class during National Walkout Day. The reports clarify that while he did not participate in the protest, he was required to be supervised, which led to his suspension. This aligns with school policies on student conduct during organized events. FAKE [Uncaptioned image] Ben Affleck & Lindsay Shookus: Planning a Wedding Even Though He’s Still Married?! It’s been almost a year since the world first learned that Ben Affleck and Lindsey Shookus are dating, and these days, the couple is reportedly ready to take the next step together.Unfortunately, there are some sizable barriers blocking their path. Verdict: UNCERTAIN; Confidence: 70 Justification: The claim about Ben Affleck and Lindsay Shookus planning a wedding is based on reports from various sources, but there is no definitive confirmation from high-credibility outlets. While some sources suggest they are in a serious relationship, the ongoing legal marriage of Affleck to Jennifer Garner complicates the situation. Additionally, there are conflicting reports regarding Affleck’s readiness to propose, such as Us Magazine (usmagazine.com) and E! Online (eonline.com), leading to uncertainty about the claim’s validity. FAKE [Uncaptioned image] Peter Strzok, FBI Agent in Texting Scandal, Will Testify Before Congress Peter Strzok, the FBI agent who sent anti-Trump messages to a colleague while leading the investigation of Hillary Clinton’s emails and Russia’s involvement in the Trump campaign, has said he is willing to testify before the House Judiciary Committee, or any other committee that wants to speak with him… Verdict: REAL; Confidence: 90 Justification: The article accurately describes Peter Strzok’s willingness to testify before Congress regarding his involvement in the investigations into Hillary Clinton’s emails and Russian interference in the Trump campaign. This is supported by multiple high-credibility sources, including reports from CBC (cbc.ca), PBS (pbs.org), and Vox (vox.com), confirming Strzok’s testimony and the context of his anti-Trump messages. The details align with the findings of the DOJ inspector general’s report, which found no evidence that Strzok’s personal views influenced specific investigative decisions. FAKE
Appendix B Qualitative Analysis
To complement the quantitative evaluation, we conducted a qualitative study to analyze how AMPEND-LS performs on individual news instances. Importantly, the examples considered here include cases drawn from outside the benchmark datasets, demonstrating the framework’s ability to generalize to real-world scenarios. The analysis highlights not only correctly classified instances but also misclassifications, thereby providing a balanced perspective on the system’s interpretability and limitations.
B.1. Correctly Classified Cases
Correct classifications occur when the model’s predictions align with the ground truth through effective evidence retrieval and structured reasoning.
Verifiable factual reporting supported by authoritative sources: In the case of the Air India Flight 171 crash (June 12, 2025), AMPEND-LS correctly predicted REAL, corroborating the claim with multiple reliable sources, including India’s Aircraft Accident Investigation Bureau and aviation-focused outlets. The LLM justification emphasized consistent technical details (e.g., fuel control switches moving to CUTOFF) and confirmed casualties, reflecting disciplined evidence selection and precise grounding in domain-specific facts.
Speculative or sensational claims lacking credible confirmation: For the Cristiano Ronaldo transfer rumor to Inter Miami, the model correctly classified the article as FAKE, citing Ronaldo’s ongoing contract with Al Nassr and the absence of credible confirmation. The justification explicitly contrasted Messi’s verified presence at Inter Miami with the unsupported rumor of Ronaldo joining, illustrating the framework’s ability to detect fabricated narratives despite their surface plausibility.
These cases demonstrate that the LLM-generated justifications are not only concise and auditable but also grounded in verifiable evidence, reinforcing transparency and user trust.
B.2. Misclassified Cases
Despite strong overall performance, some cases revealed inherent challenges where the model’s reasoning diverged from dataset labels.
Context-sensitive claims: The Ohio student suspension during National Walkout Day was predicted as REAL with high confidence because outlets like Snopes and PolitiFact confirmed the suspension event. However, the dataset labeled it FAKE, since the suspension was linked to procedural supervision rules rather than political protest. This case highlights the difficulty of handling claims where factual correctness diverges from contextual intent.
Conflicting but credible evidence: In the Ben Affleck–Lindsay Shookus wedding rumor, AMPEND-LS output UNCERTAIN, reflecting the mixed evidence across entertainment outlets. While the dataset labeled it FAKE, the model’s cautious stance demonstrates its ability to avoid overcommitment in ambiguous settings, though benchmark scoring penalizes such conservative predictions.
Framing effects versus factual content: For Peter Strzok’s congressional testimony, the model predicted REAL, supported by reputable outlets such as CBC, PBS, and Vox. Yet the dataset labeled it FAKE, likely due to framing or rhetorical exaggeration in the original claim. Here, AMPEND-LS prioritized factual accuracy over political tone, revealing tension between annotation policy and evidence-based reasoning.
Discussion: Taken together, the qualitative study illustrates that AMPEND-LS’s strengths lie in its ability to validate factual claims through evidence-grounded reasoning and to detect fabricated or speculative narratives. Misclassifications, however, reveal limitations in handling (i) context-sensitive misinformation where truth and intent diverge, (ii) ambiguity arising from credible yet conflicting sources, and (iii) discrepancies between dataset labeling criteria and factual verification. Crucially, the LLM justifications in both correct and incorrect cases remain coherent, auditable, and evidence-aware, underscoring that errors stem from reasonable interpretations rather than hallucination or arbitrary misjudgment.
Overall, the qualitative analysis confirms that AMPEND-LS not only improves predictive performance but also provides transparent reasoning, making its decisions interpretable to both users and researchers. These insights further motivate future refinements, such as context-aware reasoning, explicit uncertainty modeling, and dataset designs that better capture rhetorical framing alongside factual accuracy.
Table 8. Cross-domain analysis
Training Data Test Data Method AccAcc $F1$ PrePre RecRec PolitiFact GossipCop MGCA 49.79 49.65 49.67 49.65 BREAK 55.43 55.08 59.87 51.00 AMPEND-LS 85.33 85.43 85.53 85.33 MMCoVaR BREAK 60.48 60.46 60.13 60.79 MGCA 55.88 52.52 52.58 52.46 AMPEND-LS 76.25 75.89 81.83 70.68 GossipCop PolitiFact MGCA 43.41 43.36 41.89 44.94 BREAK 56.97 56.92 58.35 55.55 AMPEND-LS 82.78 82.65 84.67 80.73 MMCoVaR MGCA 50.18 49.66 49.56 49.76 BREAK 61.12 57.29 58.27 56.34 AMPEND-LS 80.58 80.21 79.85 80.58 MMCoVaR PolitiFact MGCA 41.94 43.92 43.81 44.04 BREAK 71.27 68.01 69.79 66.32 AMPEND-LS 83.09 84.46 85.87 83.09 GossipCop BREAK 53.40 53.45 53.51 53.40 MGCA 53.53 53.48 53.74 53.22 AMPEND-LS 74.21 73.95 74.22 73.68 PolitiFact + GossipCop MMCoVaR MGCA 62.37 60.02 61.49 58.62 BREAK 59.45 54.71 58.77 51.17 AMPEND-LS 81.57 82.85 84.17 81.57 PolitiFact + MMCoVaR GossipCop BREAK 61.60 61.63 61.66 61.60 MGCA 64.28 64.58 63.58 65.61 AMPEND-LS 71.43 71.06 71.03 71.10 GossipCop + MMCoVaR PolitiFact BREAK 68.45 65.56 68.43 62.92 MGCA 70.56 70.60 77.32 64.95 AMPEND-LS 79.41 79.01 84.40 74.26 - PolitiFact AMPEND 88.54 88.25 90.26 86.32 GossipCop 92.68 92.57 93.04 92.10 MMCoVaR 91.27 89.19 92.11 86.45
Appendix C Cross-domain Analysis
To rigorously assess the domain generalization capability of our framework, we evaluate the framework across multiple datasets that capture diverse misinformation domains. Table 8 reports comparisons against two strong baselines: MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) and BREAK (yin2025break). MGCA represents state-of-the-art multimodal approaches, aligning textual and visual signals via cross-modal contrastive learning, while BREAK exemplifies transformer-based domain adaptation through structure- and attention-aware mechanisms. These baselines were selected not only for their competitiveness in prior literature but also because they consistently performed well in our experiments, often ranking as the second-best methods across settings. This makes them appropriate and challenging comparators for evaluating the robustness of AMPEND-LS under distribution shift. The evaluation spans three heterogeneous datasets: PolitiFact (political claims), GossipCop (entertainment news), and MMCoVaR (COVID-19 health misinformation), which together reflect the topical diversity, stylistic variation, and domain heterogeneity characteristic of real-world misinformation.
[…]
(iv) AMPEND vs. AMPEND-LS: Table 8 also provides a comparative view of domain generalizability between AMPEND and AMPEND-LS. AMPEND, which directly applied the full multi-persona reasoning pipeline without distillation, exhibited stronger cross-domain generalization than AMPEND-LS. This was expected, as AMPEND leveraged rich LLM-based reasoning and iterative evidence aggregation without compression, allowing it to adapt more effectively to unseen distributions. However, when training and testing domains coincided, AMPEND-LS achieved higher performance than AMPEND, since the SLM distillation captured and reinforced domain-specific patterns while retaining evidence-grounded signals. Importantly, despite a slight degradation relative to AMPEND in true cross-domain transfer, AMPEND-LS still demonstrated substantially higher generalizability than competitive baselines such as MGCA and BREAK.
Overall, these results establish that AMPEND-LS achieves domain-agnostic robustness, balanced precision–recall, and scalable efficiency, outperforming competitive baselines under both single-source transfer and multi-source training settings. At the same time, AMPEND serves as a reasoning backbone for maximal cross-domain generalization, while AMPEND-LS extends this capability into a practical, efficient framework suitable for real-world misinformation detection.